OpenStudy (anonymous):

Solve the following integral: ln (x+3) ---------- x + 3

OpenStudy (anonymous):

Would I let (x+3) = u?

ganeshie8 (ganeshie8):

that may work... better is to substitute : ln(x+3) = u

OpenStudy (anonymous):

Yes but then I am unsure of what to do with the bottom.

ganeshie8 (ganeshie8):

ln(x+3) = u differentiate both sides 1 ----- dx = du x+3

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

Yep got to that step

OpenStudy (anonymous):

So then I would have u 1 ---- -------- X+3 X + 3

ganeshie8 (ganeshie8):

\(\large \mathbb{\int \frac{\ln(x+3)}{x+3}dx}\) substitute \(\large \mathbb{\ln(x+3) = u}\) differentiate both sides \(\large \mathbb{\frac{1}{x+3} dx = du}\) so the integral becomes : \(\large \mathbb{\int u~ du }\)

ganeshie8 (ganeshie8):

next evaluate the integral...

OpenStudy (anonymous):

U^2/2

OpenStudy (anonymous):

ANd sub back in

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

Ahhh the x+3 's cancel Thanks!!

ganeshie8 (ganeshie8):

np.. u wlc :)